OpenStudy (anonymous):

2, 4, 13, 49, 193, . . . Which term should be removed so that the sequence follows the rule?

OpenStudy (anonymous):

There's something special about the difference of two consecutive terms. But one of those differences doesn't have that same special thing in common. What do you notice about 3 of these numbers? 4 - 2 = 2 13 - 4 = 9 49 - 13 = 36 193 - 49 = 144